# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

### v1

@@ 4.4.3 SGNNs Accurately Infer the Source of Diffusion Cascades
We evaluated SGNNs ability to identify the origin of diffusion in networks—a central challenge in computational social science, epidemiology, and security analytics. Given a noisy, partially observed cascade, the model must infer which node initiated the spread.
@@
Notably, the SGNN does not approach 100% top-1 accuracy because the true source is masked 20% of the time—meaning that in one out of every five examples, the correct answer is effectively hidden. Yet the SGNN still achieves over 80% accuracy, implying that it sometimes identifies the true source even when that node was never observed to be impacted. While some advanced inference methods can handle unobserved sources under specific assumptions [21], SGNNs achieve this performance in a general setting without explicit source modeling—demonstrating a notable advance over heuristic or equation-based approaches.
@@ 4.5 SGNNs Enable Accurate Inference of Unobservable Parameters
Estimating the basic reproduction number ($R_{0}$)—the expected number of secondary infections generated by a single infectious individual in a fully susceptible population—is central to epidemic response and modeling [22, 23, 24]. However, $R_{0}$ is inherently unobservable, and traditional estimation methods rely on simplified assumptions often violated in practice—especially under data quality issues such as underreporting, delays, and early outbreak noise, as seen in early COVID-19.
@@
We next applied the SGNN to early real-world COVID-19 data (Feb 20–Mar 21, 2020) from U.S. states and cities. During this period, empirical estimates of $R_{0}$ were highly uncertain and often underestimated due to underreporting, inconsistent testing, and data delays. Many early methods assumed exponential growth and homogeneous mixing, neglecting asymptomatic spread and reporting lags.
The SGNN, trained solely on synthetic data, produced $R_{0}$ estimates from early COVID-19 data that were more consistent with later estimates once reporting had improved. For example, early estimates for New York City placed $R_{0}$ from 1.0-2.5 [27], yet subsequent studies using more complete data revised the estimate upward to over 5 [28, 29]. The SGNN, using only data from the early February–March period, correctly produced a high $R_{0}$ estimate for NYC (6.14), consistent with later, more complete analyses, versus 2.35 for exponential growth fitting. A full table of $R_{0}$ estimates is provided in Appendix E.
@@ 4.6 Back-to-Simulation Attribution Enables Mechanistic Interpretability
Beyond predictive performance, SGNNs uniquely enable back-to-simulation attribution, allowing real-world inputs to be mechanistically interpreted by identifying which simulated regimes the model deems most similar. When forecasting COVID-19 deaths in Michigan, for example, we embed the Michigan time series into the model’s learned representation space and retrieve the top-k simulated trajectories whose embeddings are closest by cosine similarity. These are not simply simulations with superficially similar curves, but rather those the model internally maps to similar latent dynamics—reflecting what the model has learned to consider mechanistically analogous during training.
@@
Figure 4 shows violin plots of four key parameters across the top-50 retrieved simulations: case-hospitalization rate, hospitalization fatality rate, asymptomatic probability, and population size. These retrieved values align closely with known or plausible values for COVID-19 in Michigan—capturing the approximate population size (median population 10.1m, versus 10.14 actual), hospitalization and death probabilities consistent with empirical estimates, and realistic asymptomatic proportions. This alignment confirms that the SGNN is not only making accurate forecasts, but is doing so by internally grounding its predictions in simulations that resemble real-world system dynamics.
This back-to-simulation attribution enables an unprecedented level of interpretability in neural networks. Unlike post hoc feature attribution techniques like SHAP or LIME, which often fail to yield mechanistic insight in complex systems, SGNN attribution operates at the level of process similarity. In this case, it tells us which kinds of epidemics the model believes it’s observing—offering a domain-aligned window into the model’s reasoning. In doing so, SGNNs bring scientific machine learning models closer to the interpretability standards of traditional mechanistic models, while maintaining the flexibility and performance of deep learning.
@@ 5 Discussion
SGNNs offer a new role for simulation in scientific machine learning—not as fixed constraints or predictive engines, but as generators of synthetic training data. This shift enables neural networks to learn both system dynamics and the observational processes that shape real-world data, including noise, underreporting, and delays.
By training on diverse ensembles of plausible simulations, SGNNs bridge the gap between mechanistic and empirical modeling. They combine inductive biases grounded in scientific theory with the flexibility of deep learning, without requiring a fixed governing equation, hand-tuned simulator, or large labeled dataset. This design supports robustness to structural mismatch and enables generalization across both domains and task types, including forecasting, regression, classification, and parameter inference.
Because simulated data include ground-truth labels for latent parameters, SGNNs allow direct supervision for scientific quantities that are unobservable in real settings. This sidesteps the need for fragile model-fitting pipelines and enables robust inference from partial, noisy, or indirect measurements.
Through back-to-simulation attribution, the model links real-world inputs to simulated regimes it internally deems similar—revealing which underlying dynamics it associates with observed patterns. This yields predictions that are not just accurate but interpretable to scientists, allowing downstream users to reason about model outputs in mechanistic terms.
The architecture-agnostic nature of SGNNs supports fine-tuning, transfer learning, and active simulation refinement. Rather than fitting a single model and hoping it generalizes, researchers can define a space of plausible behaviors via simulation, then train neural networks to learn from that space—capturing complexity and uncertainty from the outset.
Crucially, our findings suggest that SGNNs may be learning universal latent dynamics—abstract representations of scientific structure that generalize across regimes, even in zero-shot settings. For instance, models trained only on human-to-human transmission accurately predicted vector-borne dengue, suggesting internalization of cross-cutting principles (in this case harkening back to the earliest uses of SIR models for malaria [30]). This points toward simulation-pretrained foundation models for science: generalizable, low-supervision models with embedded scientific reasoning.
In doing so, SGNNs open a new direction for scientific machine learning—one that combines the rigor of mechanistic reasoning with the scalability of modern deep learning, extending the reach of modeling to problems previously beyond the scope of either approach alone.
@@ A.1 Infectious Disease Simulation Framework
$\alpha$ $\sim\text{Uniform}(0.3,1.0)$ Relative transmissibility of $A$ vs. $I$
@@
$N$ $\text{LogUniform}(5\times 10^{4},5\times 10^{7})$ Total population size
@@
Importation rate $\text{LogUniform}(0.0001,0.01)$ Daily chance of importation into $E$ or $I$
@@ A.3 Synthetic Data Generation for Suzuki-Miyaura Cross-Coupling
$y\sim\mathcal{N}(\hat{y},\sigma^{2}(\hat{y}))$ (4)
The variance function $\sigma^{2}(\hat{y})$ was learned using residuals from the real dataset, stratified into 40 equal-count bins across predicted yield levels. This captures the increased uncertainty at intermediate-to-high yields and tighter calibration at the extremes. All sampled yields were then rescaled to match the empirical mean and standard deviation via z-score normalization, and clipped to the physical interval $[0.001,0.999]$.
@@
Failure rate (yield $<$ 5%) 9.1%
@@ E.2 State $R_{0}$ Estimates
  • High-$R_{0}$ regions: NYC, NY, MA, CA, LA, MT, VI, and WA showed early $R_{0}$ values exceeding 4. These are consistent with retrospective studies that found substantial underreporting and rapid early spread, particularly in NYC, where early estimates underestimated true transmission by 2–3$\times$ [28, 29].

### v2

@@ 4.4.3 SGNNs Accurately Infer the Source of Diffusion Cascades
We evaluated SGNNs ability to identify the origin of diffusion in networks, a central challenge in computational social science, epidemiology, and security analytics. Given a noisy, partially observed cascade, the model must infer which node initiated the spread.
@@
The SGNN does not approach 100% top-1 accuracy because the true source is masked 20% of the time, meaning that in one out of every five examples, the correct answer is effectively hidden. Yet the SGNN still achieves over 80% accuracy, implying that it sometimes identifies the true source even when that node was never observed to be impacted. While some advanced inference methods can handle unobserved sources under specific assumptions [27], SGNNs achieve this performance in a general setting without explicit source modeling, demonstrating an advance over heuristic or equation-based approaches.
@@ 4.5 SGNNs Enable Accurate Inference of Unobservable Parameters
Estimating the basic reproduction number ($R_{0}$)—the expected number of secondary infections generated by a single infectious individual in a fully susceptible population—is central to epidemic response and modeling [28, 29, 30]. However, $R_{0}$ is inherently unobservable, and traditional estimation methods rely on simplified assumptions often violated in practice, especially under data quality issues such as underreporting, delays, and early outbreak noise, as seen in early COVID-19.
@@
We next applied the SGNN to early real-world COVID-19 data (Feb 20-Mar 21, 2020) from U.S. states and cities. During this period, empirical estimates of $R_{0}$ were highly uncertain and often underestimated due to underreporting, inconsistent testing, and data delays. Many early methods assumed exponential growth and homogeneous mixing, neglecting asymptomatic spread and reporting lags.
The SGNN, trained solely on synthetic data, produced $R_{0}$ estimates from early COVID-19 data that were more consistent with later estimates once reporting had improved. For example, early estimates for New York City placed $R_{0}$ from 1.0-2.5 [33], yet subsequent studies using more complete data revised the estimate upward to over 5 [34, 35]. The SGNN, using only data from the early February-March period, correctly produced a high $R_{0}$ estimate for NYC (6.14), consistent with later, more complete analyses, versus 2.35 for exponential growth fitting. A full table of $R_{0}$ estimates is provided in Appendix E.
@@ 4.6 Back-to-Simulation Attribution Enables Mechanistic Interpretability
Beyond predictive performance, SGNNs enable back-to-simulation attribution, allowing real-world inputs to be mechanistically interpreted by identifying which simulated regimes the model deems most similar. Let $f_{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ denote the trained SGNN encoder that maps input time series to a $d$-dimensional embedding space. For a real-world observation $x$, we retrieve simulations $\{x_{i}\}$ from the training set $\mathcal{D}_{\text{syn}}$ by computing cosine similarity between $f_{\phi}(x)$ and $f_{\phi}(x_{i})$, selecting the top-$k$ most similar embeddings.
This approach differs from direct time series comparison in two ways. First, retrieval in the fixed $d$-dimensional embedding space avoids the curse of dimensionality that would arise from comparing raw high-dimensional time series or parameter spaces. The embedding dimension remains constant regardless of forecast horizon or model complexity. Second, the retrieved simulations reflect the model’s learned similarity rather than time series pattern matching. The encoder $f_{\phi}$ has been trained to map trajectories to representations that are predictive of outcomes. Thus, simulations with similar embeddings are likely the ones the model treats as mechanistically similar for the prediction task.
@@
Figure 4 shows violin plots of four key parameters across the top-50 retrieved simulations: case-hospitalization rate, hospitalization fatality rate, asymptomatic probability, and population size. These retrieved values align closely with known or plausible values for COVID-19 in Michigan, capturing the approximate population size (median population 10.1m, versus 10.14 actual), hospitalization and death probabilities consistent with empirical estimates, and realistic asymptomatic proportions. This alignment confirms that the SGNN grounds its predictions in simulations that resemble real-world system dynamics.
Back-to-simulation attribution enables a different form of interpretability than feature attribution and parameter inference methods. Unlike SHAP or LIME [13, 14], which identify which input features influenced a prediction, attribution reveals which mechanistic processes the model believes generated the data. Unlike simulation-based inference approaches that aim to provide statistically rigorous parameter estimation, back-to-simulation attribution explains the model’s predictions by analyzing which mechanistic scenarios it considers similar to the observed data. This provides a domain-aligned window into the model’s predictive reasoning while maintaining computational efficiency.
@@ 5 Discussion
SGNNs provide a unified framework for incorporating mechanistic knowledge into machine learning through training data. By generating synthetic datasets from mechanistic simulators, this approach enables neural networks to learn both system dynamics and the observational processes that shape real-world data, including noise, underreporting, and delays.
Simulation-grounded learning bridges mechanistic and empirical modeling by embedding scientific structure through examples rather than architectural constraints or loss functions. This approach is compatible with any architecture and optimization method, supporting transfer learning, fine-tuning, and straightforward integration with existing deep learning infrastructure. Because simulations span diverse model structures and parameter regimes, SGNNs can be robust to structural mismatch between training simulations and real-world systems.
Training on simulated data provides ground-truth labels for quantities that are unobservable in real settings, like transmission rates or carrying capacities that lack direct measurements. This enables approaches like simulation-based inference (SBI), which use supervised learning for parameter inference. Our results show that various approaches are possible within this framework: from full Bayesian posterior estimation, established with SBI, to simpler methods like quantile regression, which can achieve accurate parameter inference with reduced computational complexity.
Back-to-simulation attribution links real-world inputs to training simulations the model considers similar, revealing which mechanistic regimes the model associates with observed patterns. Because each retrieved simulation has known ground-truth parameters and generative processes, users can inspect what the model believes is happening mechanistically. For example, if a model retrieves simulations with high asymptomatic transmission when shown an outbreak with few reported cases, this suggests the model interprets the pattern as indicating substantial hidden spread. Unlike feature attribution methods that identify which input variables influenced a prediction, back-to-simulation attribution explains which scientific processes the model believes are active. This provides actionable insight. If retrieved simulations consistently show specific parameter regimes or structural features, practitioners can validate these mechanistic hypotheses against domain knowledge or use them to guide intervention strategies.
Our results suggest that simulation-grounded learning can capture generalizable representations of scientific structure. Models trained only on human-to-human transmission accurately predicted vector-borne dengue dynamics, indicating that neural networks can internalize abstract principles that transfer across mechanistically distinct settings. This points toward the potential for simulation-pretrained foundation models in science, though substantial work remains to understand the scope and limits of such generalization.
Despite these strengths, SGNNs have important limitations. First, performance depends fundamentally on simulation quality: if key mechanisms are absent from the training simulators, the model cannot learn them. While training on diverse simulator ensembles provides robustness to individual model misspecification, systematic gaps across all training simulations will propagate to predictions. Second, generating comprehensive simulation corpora can be computationally expensive, particularly for high-fidelity mechanistic models. Though this is a one-time cost and inference remains fast, it may be prohibitive for extremely expensive simulators. Third, the sim-to-real gap remains a challenge. Even with realistic noise models and observational artifacts, systematic differences between synthetic and real data can hurt performance. Also, we lack theoretical guarantees about when SGNNs will generalize. While our results show promising zero-shot transfer (e.g., human-to-human models predicting vector-borne disease), we cannot predict a priori which cross-domain generalizations will succeed. Finally, back-to-simulation attribution reveals what the model believes is happening mechanistically, but that can be wrong. Retrieved simulations reflect learned associations that, while often informative, can be incorrect. Understanding when simulation-grounded learning succeeds and when it fails remains an important direction for future work.
SGNNs demonstrate that mechanistic simulations can serve as effective training data across diverse scientific domains, providing a framework for when and how this approach succeeds. By positioning simulation-grounded learning within the broader landscape of hybrid modeling approaches, we establish design principles for combining scientific theory with deep learning flexibility.
@@ A.1 Infectious Disease Simulation Framework
$\alpha$ $\sim\text{Uniform}(0.3,1.0)$ Relative transmissibility of $A$ vs. II
@@
NN $\text{LogUniform}(5\times 10^{4},5\times 10^{7})$ Total population size
@@
Importation rate $\text{LogUniform}(0.0001,0.01)$ Daily chance of importation into $E$ or II
@@ A.3 Synthetic Data Generation for Suzuki-Miyaura Cross-Coupling
$)y\sim\mathcal{N}(\hat{y},\sigma^{2}(\hat{y}$)) (5)
[…]
@@
Failure rate (yield << 5%) 9.1%
@@ E.2 State $R_{0}$ Estimates
[…]